Milar Corporation makes a product with the following standard costs:

Standard Quantity or
Hours
Standard Price or
Rate
Direct materials 2.0 pounds $ 7.00 per pound
Direct labor 0.5 hours $ 13.00 per hour
Variable overhead 0.5 hours $ 7.00 per hour

In January the company produced 4,500 units using 10,110 pounds of the direct material and 2,090 direct labor-hours. During the month, the company purchased 10,680 pounds of the direct material at a cost of $76,560. The actual direct labor cost was $38,257 and the actual variable overhead cost was $11,958.

The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased.

The materials price variance for January is:

Multiple Choice

  • $1,660 F

  • $1,800 U

  • $1,660 U

  • $1,800 F
Answer #1

Materials Price Variance = (SP - AP) * AQ

= (7 - 76560/10680) * 10680

= (7 - 7.17) * 10680

= 1,800 U

Option B